The displacement vectors A and B, when added together, give the resultant vector R, so that R = A + B. Use the data in the drawing and the fact that φ = 26° to find the magnitude R of the resultant vector and the angle θ that it makes with the +x axis.



1
Expert's answer
2021-01-31T18:58:02-0500
"R=\sqrt{6^2+8^2+2(6)(8)\cos{(90+26)}}=7.6\ m"

"8\cos{26}=7.6 \cosθ\\θ=\arccos0.9461=19°"